A is the point (2,3) and B is the point (4,9)
(a) Find the slope of the line segment [AB]
(b)Find the slope of a line perpendicular to line segment [AB]
(c)The line 2x+by-12=0 is perpendicular to the line segment [AB].What is the value of b?

Answer by checkley77(12844)

(2,3)(4,9)
Slope=(y2-y1)/(x2-x1)
Slope=(9-3)/(4-2)
Slope=6/2=3 ans.
--------------------
A perpendicular line will have a slope=-1/3.
-----------------------------------------------
2x+by-12=0
by=-2x+12
y=-2x/b+12/b [b=6]
y=-2x/6+12/6
y=-x/3+2